Shift report 20010801 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Aug 01 2001 - 23:52:51 EDT

  • Next message: Apache: "Shift report 20010802 00:00-08:00"

    ------------------------------------------------------------
    Date: Wed Aug  1 23:52:50 2001, Shift: 16:00-24:00
    Supervisor: P. C., with: K. H.
    Log book#6 pages 143 - 146
    
    Shift Summary:
    ------------------------------------------------------------
    No beam during entire shift.
    
    Machine status:
    First part was the end of scheduled maintenance.
    
    During the rest of the time, Status was machine 
    development.
    
    At the moment, Gold is in AGS, RHIC is swept and preparations
    for beam operation continue.
    
    Experimental status:
    FFS was moved to 6 deg
    Current positions:
    FFS 6 deg
    BFS 6 deg
    MRS 90 deg
    
    All magnets turned on to 1/5 max field
    D1 676 (2.5 kG) pol A
    D2 555 (3.4 kG) pol A
    D3 588 (2.4 kG) pol A
    D4 534 (2.4 kG) pol A
    D5 352 (2.1 kG) pol A (1/7 max field)
    
    These are the conditions to start Second Goal step 1
    
    For First Goal step 3, we have about 333k events, but it
    should be noted that T1 was not working during the whole
    time and that part might well have to be done over.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Aug 01 2001 - 23:53:35 EDT